OpenStudy (anonymous):

simplify 17 . 24 divide (17 . 24) so from what i understand you start with the 17 . 24.. tay sart left to right ?

OpenStudy (anonymous):

you start*

OpenStudy (anonymous):

What is that period supposed to be?

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

x/x=1

OpenStudy (anonymous):

=1

OpenStudy (anonymous):

so the problem looks like this: \[17 \times 24/(17 \times 24)\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Does the problem you are having a question about look like the one I just posted?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

Well it's a good habit to use order of operations when solving. In this case parenthesis then multiplication/division, but since the problem is in the form x/x and x does not equal zero the the answer is 1.

OpenStudy (anonymous):

then it's r8..ok ?